A jogger runs into the country side at a rate of 10 mph. He returns along the same route at 6mph. If the total trip took 1 hour 36 minutes, how far did he jog?
Thank you!

Answer by ankor@dixie-net.com(22740)

A jogger runs into the country side at a rate of 10 mph. He returns along the same route at 6 mph. If the total trip took 1 hour 36 minutes, how far did he jog?
:
Let d = distance one way: Change 36 min to hrs: 36/60 = .6 hrs
Time = dist/speed
:
Write a time equation, find d:
Time out + time back = 1.6 hrs
d/10 + d/6 = 1.6
:
Get rid of the denominator, mult eq by 30, resulting in:
3d + 5d = 1.6(30)
8d = 48
d = 48/8
d = 6 mi one way. Total he jogged = 12 mi
:
Check: 6/10 + 6/6 = 1.6